# The algebra $\mathcal{D}(W)$ via strong Darboux transformations

### v1

@@ 5.1. Example I
Let us prove that there are no operators of order 2 or 4 in $\mathcal{Z}(W)$. If such operators existed, its leading coefficient is $\alpha I$, for some $\alpha\in\mathbb{C}$. Then it would be a linear combination of $D_{1}^{2}$, $D_{1}$, and $I$. But, $\beta_{1}D_{1}^{2}+\beta_{2}D_{1}+\mu I$ does not commute with $D_{2}$ for any $\beta_{1},\beta_{2}\neq 0$.
Recall that all differential operator $D\in\mathcal{D}(W)$ has even order and let $D=\sum_{j=0}^{2m}\partial^{j}F_{j}\in\mathcal{Z}(W)$. By Proposition 5.9 we have that $F_{2m}=\alpha I$ for some $\alpha\in\mathbb{C}$. We proceed by induction on $m$. For $m=3$, we have that $D$ is of order $6$ and $F_{6}=\alpha I$ for some $\alpha\in\mathbb{C}$. Then, $D-\alpha Z_{1}$ belongs to $\mathcal{Z}(W)$ and it is of order less than $6$. Thus, $D=\alpha Z_{1}+kI$ for some $k\in\mathbb{C}$.
@@ 5.2. Example II
As we mentioned before, the algebra $\mathcal{D}(W)$ associated with this weight has already been studied by J. Tirao in [26], a comprehensive and technical paper. Here, we provide an alternative and shorter method to compute this algebra, as a consequence of the fact that the weight $W$ is a Darboux transformation of $w$.
@@
The weight $W(x)$ is a strong Darboux transformation of $w(x)=e^{-x^{2}}I$.
@@
$D=\begin{pmatrix}-\delta+2+\frac{4}{a^{2}}&&0\\ 0&&-\delta+\frac{4}{a^{2}}\end{pmatrix}\in\mathcal{D}(w)$
@@
(46) $\begin{split}\mathcal{V}&=\partial\begin{pmatrix}0&&1\\ 1&&-ax\end{pmatrix}-\frac{2}{a}I,\\ \mathcal{N}=W\mathcal{V}^{\ast}w^{-1}&=\partial\begin{pmatrix}-ax&&-1\\ -1&&0\end{pmatrix}+\begin{pmatrix}-a-\frac{2}{a}&&0\\ 0&&-\frac{2}{a}\end{pmatrix}.\end{split}$
By integration by parts, we easily obtain that $\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{w}$, for all $P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C})[x]$. Thus, by Proposition 3.4, the statement holds. ∎
From Theorem 3.9 we can map $w$-symmetric operators into $W$-symmetric operators. Thus we have the following (second-order) $W$-symmetric differential operators in the algebra $\mathcal{D}(W)$.
@@
where $E_{1}=\begin{pmatrix}0&&1\\ 1&&0\end{pmatrix}$, $E_{2}=\begin{pmatrix}0&&0\\ 0&&1\end{pmatrix}$, $E_{3}=\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}$ and $E_{4}=i\begin{pmatrix}0&&1\\ -1&&0\end{pmatrix}$, and $\mathcal{V}$, $\mathcal{N}$ are defined in (46). We also introduce the differential operator $D_{5}$, whose leading coefficient is a scalar matrix
@@
The weight $W$ is a strong Darboux transformation of $e^{-x^{2}}I$, thus by Corollary 3.11, we have that all differential operators in $\mathcal{D}(W)$ are of even order. If $D=\sum_{j=0}^{2n}\partial^{j}F_{j}\in\mathcal{S}(W)$ then $\mathcal{V}D\mathcal{N}\in\mathcal{S}(w)$, and thus, from Proposition 5.2, the leading coefficient of $D$ is of the form
@@ 6. Beyond the Matrix Bochner Problem: New weight matrices
Throughout this paper, we have illustrated some profound relationships that exist between the algebras of differential operators associated with weight matrices that are Darboux transformations of each other. In this section, we introduce two new families of irreducible Hermite-type weight matrices, of size $2\times 2$. Both examples can be obtained as Darboux transformation of a direct sum of classical weights $w(x)=e^{-x^{2}}I$. However, they are not solutions to the Matrix Bochner Problem, meaning that their algebra $\mathcal{D}(W)$ does not contain second-order differential operators.
@@
The weight $W$ can be obtained through a Darboux transformation of $w(x)=e^{-x^{2}}I$, a direct sum of classical scalar Hermite weights. Let $\delta=\partial^{2}-2\partial x$ be the classical Hermite second-order differential operator.
The following eighth-order $w$-symmetric differential operator
@@
belongs to the algebra $\mathcal{D}(w)$ and it can be factored as $\mathcal{D}=\mathcal{V}\mathcal{N}$, where $\mathcal{N}=W\mathcal{V}^{\ast}w^{-1}$ and
@@
$\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{w},\quad\text{ for all $P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C})[x]$}.$
In this way, by Proposition 3.4, we obtain that the weight $W$, given in (48), is a Darboux transformation of $w(x)=e^{-x^{2}}I$. As a consequence, $Q_{n}(x)=H_{n}(x)\cdot\mathcal{V}$ is a sequence of orthogonal polynomials for $W(x)$.
@@
is a strong Darboux transformation of $w(x)=e^{-x^{2}}I$ by factoring the eighth-order $w$-symmetric differential operator
$\mathcal{D}=\begin{pmatrix}{\delta}^{4}-2{\delta}^{3}+{\frac{4\left(-{a}^{2}+8\right){\delta}^{2}}{{a}^{2}}}+{\frac{8\left({a}^{4}-20\,{a}^{2}-8\,{b}^{2}\right)\delta}{{a}^{4}}}+{\frac{64(3\,{a}^{2}+2\,{b}^{2}+4)}{{a}^{4}}}&-{\frac{24\,{a}^{2}{\delta}^{2}-48\,{a}^{2}\delta-128}{{a}^{3}}}\\ -{\frac{24\,{a}^{2}{\delta}^{2}-48\,{a}^{2}\delta-128}{{a}^{3}}}&{\delta}^{4}-6\,{\delta}^{3}+{\frac{8\left({a}^{2}+4\right){\delta}^{2}}{{a}^{2}}}+{\frac{32\left({a}^{2}-2\,{b}^{2}\right)\delta}{{a}^{4}}}+\frac{256}{{a}^{4}}\end{pmatrix}\in\mathcal{D}(w)$
with strong Darboux transformer from $w$ to $W$ given by

### v2

@@ 5.1. Example I
Let us prove that there are no operators of order 2 or 4 in $\mathcal{Z}(W)$. If such an operator existed, its leading coefficient would be $\alpha I$, for some $\alpha\in\mathbb{C}$. Then it would be a linear combination of $D_{1}^{2}$, $D_{1}$, and $I$. But, $\beta_{1}D_{1}^{2}+\beta_{2}D_{1}+\mu I$ does not commute with $D_{2}$ for any $\beta_{1},\beta_{2}\neq 0$.
Recall that every differential operator $D\in\mathcal{D}(W)$ has even order and let $D=\sum_{j=0}^{2m}\partial^{j}F_{j}\in\mathcal{Z}(W)$. By Proposition 5.9 we have that $F_{2m}=\alpha I$ for some $\alpha\in\mathbb{C}$. We proceed by induction on $m$. For $m=3$, we have that $D$ is of order $6$ and $F_{6}=\alpha I$ for some $\alpha\in\mathbb{C}$. Then, $D-\alpha Z_{1}$ belongs to $\mathcal{Z}(W)$ and it is of order less than $6$. Thus, $D=\alpha Z_{1}+kI$ for some $k\in\mathbb{C}$.
@@ 5.2. Example II
As we mentioned before, the algebra $\mathcal{D}(W)$ associated with this weight has already been studied by J. Tirao in [26], a comprehensive and technical paper. Here, we provide an alternative and shorter method to compute this algebra, as a consequence of the fact that the weight $W$ is a Darboux transformation of $\widetilde{W}$.
@@
The weight $W(x)$ is a strong Darboux transformation of $\widetilde{W}(x)=e^{-x^{2}}I$.
@@
$D=\begin{pmatrix}-\delta+2+\frac{4}{a^{2}}&&0\\ 0&&-\delta+\frac{4}{a^{2}}\end{pmatrix}\in\mathcal{D}(\widetilde{W})$
@@
(22) $\begin{split}\mathcal{V}&=\partial\begin{pmatrix}0&&1\\ 1&&-ax\end{pmatrix}-\frac{2}{a}I,\\ \mathcal{N}=W\mathcal{V}^{\ast}\widetilde{W}^{-1}&=\partial\begin{pmatrix}-ax&&-1\\ -1&&0\end{pmatrix}+\begin{pmatrix}-a-\frac{2}{a}&&0\\ 0&&-\frac{2}{a}\end{pmatrix}.\end{split}$
By integration by parts, we easily obtain that $\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{\widetilde{W}}$, for all $P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C}[x])$. Thus, by Proposition 3.4, the statement holds. ∎
From Theorem 3.9 we can map $\widetilde{W}$-symmetric operators into $W$-symmetric operators. Thus we have the following (second-order) $W$-symmetric differential operators in the algebra $\mathcal{D}(W)$.
@@
where $E_{1}=\begin{pmatrix}0&&1\\ 1&&0\end{pmatrix}$, $E_{2}=\begin{pmatrix}0&&0\\ 0&&1\end{pmatrix}$, $E_{3}=\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}$ and $E_{4}=i\begin{pmatrix}0&&1\\ -1&&0\end{pmatrix}$, and $\mathcal{V}$, $\mathcal{N}$ are defined in (22). We also introduce the differential operator $D_{5}=-\mathcal{N}\mathcal{V}$, whose leading coefficient is a scalar matrix
@@
The weight $W$ is a strong Darboux transformation of $e^{-x^{2}}I$, thus by Corollary 3.11, we have that all differential operators in $\mathcal{D}(W)$ are of even order. If $D=\sum_{j=0}^{2n}\partial^{j}F_{j}\in\mathcal{S}(W)$ then $\mathcal{V}D\mathcal{N}\in\mathcal{S}(\widetilde{W})$, and thus, from Proposition 5.2, the leading coefficient of $D$ is of the form
@@ 6. Beyond the Matrix Bochner Problem: New weight matrices
Throughout this paper, we have illustrated some profound relationships that exist between the algebras of differential operators associated with weight matrices that are Darboux transformations of each other. In this section, we introduce two new families of irreducible Hermite-type weight matrices, of size $2\times 2$. Both examples can be obtained as Darboux transformations of a direct sum of classical weights $\widetilde{W}(x)=e^{-x^{2}}I$. However, they are not solutions to the Matrix Bochner Problem, meaning that their algebra $\mathcal{D}(W)$ does not contain second-order differential operators.
@@
The weight $W$ can be obtained through a Darboux transformation of $\widetilde{W}(x)=e^{-x^{2}}I$, a direct sum of classical scalar Hermite weights. Let $\delta=\partial^{2}-2\partial x$ be the classical Hermite second-order differential operator.
The following eighth-order $\widetilde{W}$-symmetric differential operator
@@
belongs to the algebra $\mathcal{D}(\widetilde{W})$ and it can be factored as $\mathcal{D}=\mathcal{V}\mathcal{N}$, where $\mathcal{N}=W\mathcal{V}^{\ast}\widetilde{W}^{-1}$ and
@@
$\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{\widetilde{W}},\quad\text{ for all $P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C}[x])$}.$
In this way, by Proposition 3.4, we obtain that the weight $W$, given in (24), is a Darboux transformation of $\widetilde{W}(x)=e^{-x^{2}}I$. As a consequence, $Q_{n}(x)=H_{n}(x)\cdot\mathcal{V}$ is a sequence of orthogonal polynomials for $W(x)$.
@@
is a strong Darboux transformation of $\widetilde{W}(x)=e^{-x^{2}}I$ by factoring the eighth-order $\widetilde{W}$-symmetric differential operator
$\mathcal{D}=\begin{pmatrix}{\delta}^{4}-2{\delta}^{3}+{\frac{4\left(-{a}^{2}+8\right){\delta}^{2}}{{a}^{2}}}+{\frac{8\left({a}^{4}-20\,{a}^{2}-8\,{b}^{2}\right)\delta}{{a}^{4}}}+{\frac{64(3\,{a}^{2}+2\,{b}^{2}+4)}{{a}^{4}}}&-{\frac{24\,{a}^{2}{\delta}^{2}-48\,{a}^{2}\delta-128}{{a}^{3}}}\\ -{\frac{24\,{a}^{2}{\delta}^{2}-48\,{a}^{2}\delta-128}{{a}^{3}}}&{\delta}^{4}-6\,{\delta}^{3}+{\frac{8\left({a}^{2}+4\right){\delta}^{2}}{{a}^{2}}}+{\frac{32\left({a}^{2}-2\,{b}^{2}\right)\delta}{{a}^{4}}}+\frac{256}{{a}^{4}}\end{pmatrix}\in\mathcal{D}(\widetilde{W})$
with strong Darboux transformer from $\widetilde{W}$ to $W$ given by